OpenStudy (anonymous):

How much money, invested at an interest rate of r % per year compounded continuously, will amount to A dollars after t years? (Round your answer to the nearest cent.) A = 17,000, r = 4.3, t = 3

OpenStudy (anonymous):

A= Pe^rt right..?

OpenStudy (anonymous):

but how will i round it to the nearest cent..?

OpenStudy (dumbcow):

\[A = Pe^{rt}\] solving for P \[P = \frac{A}{e^{rt}}\] yeah, plug in the values

OpenStudy (anonymous):

ow.. derive formula..

OpenStudy (anonymous):

so it is 14,942.55741 how will i round this to the nearest cent..?

OpenStudy (dumbcow):

round to 2 decimals

OpenStudy (anonymous):

ahhh.. tnx.. xDDDD

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

Find the remainder r if the polynomial 4x^100 + 4x^85 − 4x^38 + 2x^17 − 6 divide by x+1 how can i find..?? will i input in the synthetic divison.. 4 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 ..? up to the last to get the remainder..?

OpenStudy (anonymous):

or there is a shortcut..?

OpenStudy (dumbcow):

lol synthetic division is the short cut

OpenStudy (dumbcow):

crazy problem hope you have a lot of room

OpenStudy (anonymous):

yeah.. but 4x^100.. is so long to do..?

OpenStudy (dumbcow):

hold on, you just need the remainder right?

OpenStudy (anonymous):

yep.. just that crazy remainder.. hahah

OpenStudy (dumbcow):

then there is something called the remainder theorem :) f(c) = remainder of dividing polynomial by (x-c) so just plug in x = -1 to get remainder

OpenStudy (anonymous):

huh.. wait didnt get it..? i will just substitute it..?

OpenStudy (dumbcow):

yeah forget dividing , just evaluate at x=-1

OpenStudy (anonymous):

hwahha.. and then..? perform the equation after substituting..?

OpenStudy (dumbcow):

i could show the proof but im sure its in your textbook or notes somewhere

OpenStudy (dumbcow):

what do you get after substituting?

OpenStudy (anonymous):

the good news is maybe it is in somewhere in my notes or textbooks but the bad news is.. i forgot it in my locker at school... haha

OpenStudy (anonymous):

-12.. hwahah.. is that it..?

OpenStudy (dumbcow):

no worries http://en.wikipedia.org/wiki/Polynomial_remainder_theorem#Proof

OpenStudy (dumbcow):

yes its -12

OpenStudy (anonymous):

wow awesome.. thats cool.. its my first time to know that u can use the remainder thoerem at this one.. haha.. i got some tips now.. hwahah

OpenStudy (dumbcow):

:)

OpenStudy (anonymous):

last.. forgot to use the long division.. is synthetic only applicable if the x is single variable and has 1 exponent..? Find the quotient and remainder if f(x) is divided by p(x). f(x) = 3x^3 − 5x^2 − 4x − 12; p(x) = 2x^2 + x

OpenStudy (dumbcow):

right it has to be in the form (x-c) to use synthetic division so any x^2 will not work

OpenStudy (dumbcow):

hey if you need help with long division, here is a video https://www.khanacademy.org/math/algebra2/polynomial_and_rational/dividing_polynomials/v/dividing-polynomials-with-remainders its harder to help with long division using text

OpenStudy (anonymous):

yeah.. i guess haha.. because i can recall some steps but not all.. hahha

OpenStudy (anonymous):

tnx for the link..